OpenStudy (anonymous):

The equation, y = 10(1.0605)x , represents the winning prize money, in thousands, for first place in the Indianapolis 500 auto race. x represents the number of years after 1911, the first year. What is the amount of prize money in 1961? Use x = 50, as it is 50 years after 1911. Round to the nearest whole number and add three zeros to the end. For instance, if you get an answer like 142.8, this rounds to 143, and add 3 zeros to get $143,000. A. $189,000 B. $530,000 C. $80,000

OpenStudy (anonymous):

hi, do you need help with this?

hero (hero):

Don't just stare at it. Provide a solution.

OpenStudy (anonymous):

A solution.

OpenStudy (anonymous):

if you're not part of the solution you're part of the precipitate

OpenStudy (anonymous):

its B) $530,000 y = 10 *(1.0605)* 50 y = 530.25 round off to the nearest number and you get 530 add 3 zeros and you get $530,000

OpenStudy (anonymous):

B. y=10*1.0605*50 y=500*1.0605 y=530.25 rounded off to y=530 adding 3 zeros Ans= $530000